OpenStudy (anonymous):

prove secx+tanx+k then prove that sinx=k*k-1\k*k+1

OpenStudy (mr.math):

That doesn't make sense!

OpenStudy (anonymous):

May be it makes sense to in some higher dimension?

OpenStudy (mr.math):

Could you rewrite the question again, and make sure you write it correctly?!

OpenStudy (anonymous):

Yeah it might be when secx+tanx=k prove sinx = above said expression in k. Proof goes on here:- Rewriting secx and tanx we get (1+sinx)/cosx = k now multiply and divide by (1-sinx) Now we get cosx/(1-sinx)=k Implies cosx=k(1-sinx) Substitute this value of cosx in above we get (1+sinx)/k(1-sinx) = k Implies (1+sinx)/(1-sinx) = k^2 (1+sinx)=k^2(1-sinx) 1+sinx=k^2-k^2sinx (k^2+1)sinx=k^2-1 sinx=(k^2-1)/(k^2+1) Hence Proved ur result.... Any doubts in this Please do comment